OpenStudy (anonymous):

8. Solve. 3/2x – 4 = 16

OpenStudy (anonymous):

3/2x=20 x=40/3

jhonyy9 (jhonyy9):

3/2x=20 3=40x x=3/40

jhonyy9 (jhonyy9):

- ATTENTION Remainder !